Allen Air Lines must liquidate some equipment that is being replaced. The equipment originally cost $22.4 million, of which 80% has been depreciated. The used equipment can be sold today for $5.6 million, and its tax rate is 40%. What is the equipment's after-tax net salvage value? Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000. Round your answer to the nearest dollar.
Answer #1

Book value as on date of sale=cost-Accumulated Depreciation

=22.4*(1-0.8)=$4.48 million

Hence gain on sale=5.6-4.48=$1.12 million

Hence after-tax net salvage value=Sale proceeds-(Tax rate*gain on sale)

=5.6-(0.4*1.12) million

=$5,152,000